40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Dec 26 17:45:21 2014, Steve, Update, VAC, vac pressure rose to 1.3 mTorr noN2.pngbacktoVacNormal.png
    Reply  Sun Dec 28 07:29:16 2014, Steve, Update, VAC, vacuum is normal 3d.png
    Reply  Tue Feb 3 15:20:13 2015, Steve, Koji, Update, VAC, Pneumatic pressure is being read again N2pressureReadingIsBack.png
    Reply  Sun Mar 22 18:42:40 2015, Steve, Update, VAC, vac pressure rose to 1.2mTorr NoN2.pngPressureReadingWorks.png
       Reply  Mon Mar 23 10:36:55 2015, ericq, Update, VAC, Pressure watch script 
          Reply  Fri Apr 24 17:30:32 2015, Jenne, Update, VAC, Pressure watch script broken 
             Reply  Sat Apr 25 18:50:47 2015, ericq, Update, VAC, Pressure watch script broken 
                Reply  Sat Apr 25 22:17:49 2015, rana, Update, CDS, MXstream restart script working (beta) 
                   Reply  Tue May 19 17:03:34 2015, Koji, Update, CDS, MXstream restart script working (beta) 
                      Reply  Tue May 19 18:38:33 2015, rana, Update, CDS, MXstream restart script working (beta) 
          Reply  Mon Apr 27 14:37:55 2015, Steve, Update, VAC, N2 pneumatic pressure watch  
             Reply  Thu Apr 30 16:30:25 2015, Steve, Update, VAC, N2 pneumatic pressure watch set up 
          Reply  Tue Jan 31 14:14:29 2017, rana, Update, VAC, Pressure watch script 
       Reply  Sun May 15 19:48:08 2016, Steve, Update, VAC, run out of N2 noN2.png
          Reply  Tue Jun 14 10:13:34 2016, Steve, Update, VAC, temporary N2 supply line afterN2work.png
             Reply  Fri Jun 17 15:22:31 2016, gautam, Update, VAC, N2 supply line restored after retiling c0vac_06172016.png
Message ID: 12195     Entry time: Fri Jun 17 15:22:31 2016     In reply to: 12174
Author: gautam 
Type: Update 
Category: VAC 
Subject: N2 supply line restored after retiling 
Quote:

The drill room floor will be retiled Thursday, June 16. Temporary nitrogen line set up will allow emptying the hole area.

 

Ifo room entry will be through control room.

 

The retiling work has finished, Steve and I restored the N2 supply configuration to its normal state. The sequence of steps followed was:

  1. Went to the X end and closed the following valves, roughly in this order: VAEE, VAEV, VABS, VABSCI, VASV, VASE, V4, V1.
  2. Checked the RPM on the various turbo pump controllers to make sure they were in their nominal states
  3. Disconnect the electrical connections to V1, V4, V5 and VA6 - just to make sure some spurious signal doesn't unintentionally open any of these valves while we are mucking around with the N2 supply
  4. Close the valves on the N2 cylinders in the drill room. Disconnect the temporary nitrogen line (at this point, the N2 pressure to the IFO valves goes down from ~7-PSI to 0), reconnect the old supply chain, taking care that we aren't unintentionally loosening any of the Swagelock connections while unscrewing stuff
  5. Replaced one of the N2 cylinders that was running low.
  6. Reopen the cylinders, restore N2 pressure to IFO valves to ~70PSI.
  7. Do steps 1-3 in reverse: i.e. reconnect power to all valves, open them in the reverse order we closed them while monitoring the state of the various turbo pumps. 
  8. Acknowledged the error message on the C0VAC medm screen

Note: the valve isolating the RGA automatically shutoff during this work, possibly because it detected a pressure above its threshold - after checking the appropriate pressure gauges, we reopened this valve as well. 

The attached screenshot suggests that everything went as planned and that the vacuum system is back to normal...

 

 

Attachment 1: c0vac_06172016.png  24 kB  | Hide | Hide all
c0vac_06172016.png
ELOG V3.1.3-